OpenStudy (anonymous):

Solve. 2x + 4 = 8 A. x = 2 B. x = 4 C. x = 6 D. x = 8

OpenStudy (a1234):

2x + 4 = 8 Subtract 4 from each side first. 2x = 4 Divide each side by 2. 2x/2 = 4/2 x = 2

OpenStudy (anonymous):

thx

OpenStudy (whpalmer4):

In general, it works best to do the operations in the opposite order of their priority: PEMDAS P=parenthese E=exponentiation M=multiplication D=division A=addition S=subtraction So when solving, do the addition and subtraction first, then do the multiplication and division, etc. That's what you did here — you did the subtraction, then the division

OpenStudy (solomonzelman):

I had in mind to divide everything by 2 first, then the answer is right there